Simplify the complex fraction?
1 + 1/x= 30/x2

Answer by benni1013(206)

In order to properly solve you need to make the equation equal to 0:
1+1/x-30/x^2=0
Multiply each side by the common denominator, in your case, x^2:
x^2+x-30=0
Factor:
(x-6)(x+5)=0
Separate the two quantity and set each equal to zero:
x-6=0 and x+5=0
Solve:
x=6,-5